what is 15 th term of A.P whose 1 st term is 25 and 3 rd term is 31

what is 15th term of A.P whose 1st term is 25 and 3rd term is 31  

Grade:7

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given 1stterm=a=25.....(1) and 3rdterm=a+2d=31......(2)
Substituting 2 in 1 we get 25+2d=31 so 2d=6 which gives d=3
So 15thterm=a+14d=25+14(3)=25+42=67